# SANeRF-HQ: Segment Anything for NeRF in High Quality

### 03 Dec 2023 v1

@@ 4.5 Ablation Study
Mask Aggregator. Based on point prompts, we perform ablation study on the mask aggregator, comparing the intermediate results from the feature container, i.e., directly decoding the SAM features from the encoder or the feature field, to the final outputs of the mask aggregator. The quantitative results are in Table 1.
@@
Directly propagating point prompts between different views and applying SAM cannot guarantee cross-view consistency. When prompts are sparse, some 3D point prompts may be occluded at certain viewpoints, and no mask can be produced. Even when a large number of prompts are provided, SAM masks may still cover different objects across views, and this naive approach fail to utilize masks from other views to collectively refine the results. The same issue also exists in SAN [10], which distills SAM features with another neural field and later decodes the rendered features using the decoder. In addition to the consistency issue, the rendered feature maps from SAN further suffer from aliasing, and fine spatial semantics in the SAM features along object boundaries can be lost during interpolation. This can lead to less accurate segmentation results on the object boundaries, like jagged mask edges. To ensure fairness in comparison, we introduce enough point prompts to guarantee masks can be generated from every viewpoint where the target object is visible.
@@
Figure 5 illustrates the qualitative comparison results. Although SANeRF-HQ uses the potentially inconsistent segmentation from SAM as input, by aggregating multi-view information and integrating 3D geometry captured by NeRF, it can still produce a underlying 3D mask close to the ground-truth geometry, which guarantees consistent multi-view masks and usually comes with higher quality.
Ray-Pair RGB Loss. We perform ablations on the Ray-Pair RGB loss in Section 3.3.2. Table 2 shows the quantitative results, where the Ray-Pair RGB loss slightly enhances the mask quality. The visual improvement of the masks illustrated in Figure 7 is more significant. The loss helps fill the missing interior and boundaries of the masks by enforcing a local match between the similarity in labels, and the similarity in appearance.
For more ablation study, please refer to our supplementary materials.
Figure 7: Qualitative Results of Ray-Pair RGB Loss. The Ray-Pair RGB loss can help to recover local regions and make the results more solid.
@@ 5 Conclusion
In this paper, we introduced the Segment Anything for NeRF in High Quality (SANeRF-HQ) framework. By combining the strengths of the Segment Anything Model (SAM) for open-world object segmentation and NeRF for aggregating information from multiple viewpoints, SANeRF-HQ represents a significant advancement in high-quality 3D segmentation. Our method was quantitatively and qualitatively evaluated on various NeRF datasets, which demonstrates SANeRF-HQ’s advantages over previous state-of-the-art methods. Furthermore, we demonstrate the potential of extending our work to 4D dynamic NeRF object segmentation (see supplementary materials). SANeRF-HQ holds promise for contributing significantly to the evolving landscape of 3D computer vision and segmentation techniques.
@@ Appendix A Implementation Details
Appendix B More Ablation Study
SAM Backbones. We perform ablation study on different types of backbones for feature extracting and decoding, which are SAM and SAM-HQ [29].
Feature Container Method. We propose two methods to store the pre-computed features in feature container Section 3.1. The first one is to store the feature maps in a cache. The second one is to distill the features into a feature field and render the feature map during inference. We investigate the performance of these two methods on our evaluation set.
Table C.3 demonstrates our the quantitative results on the ablation study. SAM-HQ with a cache container performs the best in most cases, while SAM-HQ with a distillation container does not produce similar results as others on our evaluation set. Nevertheless, our method demonstrates advantages over other 3D segmentation methods in general. Notice that we do not choose the best results from all of them in our comparison to ensure the fairness. Considering that all the previous methods leverage SAM, we use SAM with feature distillation as our feature container in comparison.
@@ Appendix C Comparison with Instance Segmentation Methods
Table C.2: Comparison with Panoptic Lifting and Contrastive Lift on the data mentioned in their papers.
@@
Backbone Container Mip-NeRF 360 LERF LLFF 3D-FRONT Others
SAM Cache 99.3 93.6 98.9 90.3 99.5 95.8 97.6 84.8 99.6 91.1
Distillation 99.2 91.0 99.0 90.7 99.3 95.2 98.6 89.9 99.6 91.1
SAM-HQ Cache 99.4 94.4 99.2 93.1 99.6 96.8 98.9 91.8 99.5 90.1
Distillation 98.9 88.1 98.0 86.3 98.8 89.8 96.3 79.4 99.4 89.4
Table C.3: Quantitative Results on different backbones and feature containers.
@@ Appendix E More Qualitative Results
We demonstrate extra qualitative comparisons between our method and other zero-shot 3D segmentation methods as mentioned in the main paper. The results are given in Figures E.1E.2E.3E.4. Please watch the video for more qualitative results.
Figure E.1: Comparison with SA3D and ISRF on the Replica Room. Data is from the Others subset. SANeRF-HQ can maintain the object structure while excludes the background.
Figure E.2: Comparison with SA3D and ISRF on the Shoe Rack. Data is from the LERF subset. Our method can reproduce the segmentation details even with some occlusion.
Figure E.3: Comparison with SA3D and ISRF on the Hypersim. Data is from the Others subset. ISRF contains too many false positives, while SA3D cannot cover the whole object.
Figure E.4: Comparison with SA3D and ISRF on the Espresso. Data is from the LERF subset. Our method produces the most reasonable segmentation in the distant, complex setting.
ottom:-0.4ex;">Exml[LOGO]

### 06 Apr 2024 v2

@@ 4.4 Ablation Studies
Mask Aggregator. Based on point prompts, we perform ablation study on the mask aggregator, comparing the intermediate results from the feature container, i.e., directly decoding the SAM features from the encoder or the feature field, to the masks after aggregating. The quantitative results are in Table 1.
@@
Directly propagating point prompts between different views and applying SAM cannot guarantee cross-view consistency. When prompts are sparse, some 3D point prompts may be occluded at certain viewpoints and no mask can be produced. Even when a large number of prompts are provided, SAM masks may still cover different objects across views, and this naive approach fail to utilize masks from other views to collectively refine the results. The same issue also exists in SAN, which distills SAM features with another neural field and later decodes the rendered features using the decoder. In addition to the consistency issue, the rendered feature maps from SAN further suffer from aliasing, and fine spatial semantics in the SAM features along object boundaries can be lost during interpolation. This can lead to less accurate segmentation results on the object boundaries, like jagged mask edges. To ensure fairness in comparison, we introduce enough point prompts to guarantee masks can be generated from every viewpoint where the target object is visible.
@@
Figure 5 illustrates the qualitative comparison results. Although SANeRF-HQ uses the potentially inconsistent segmentation from SAM as input, by aggregating multi-view information and integrating 3D geometry captured by NeRF, it can produce a underlying 3D mask close to the ground-truth geometry. This guarantees consistent multi-view masks and usually comes with higher quality.
Ray-Pair RGB Loss. Table 3 shows the quantitative results of ablations on the Ray-Pair RGB loss in Section 3.3.2, where the Ray-Pair RGB loss slightly enhances the mask quality. We refer readers to the supplementary material for qualitative comparison.
@@
4.5 More Qualitative Results
Figure 6 illustrates the results of our methods on other segmentation tasks. With Grounding-DINO [37], our method can segment objects based on text prompts. Additionally, our method can perform automatic 3D segmentation, utilizing videos rendered by NeRF, the auto-segmentation function of SAM, and incorporating [17] into the mask decoder.
More qualitative results, comparisons, preliminary results on dynamic NeRF segmentation, and discussion on limitations can be found in the supplementary material.
Figure 6: Qualitative Results of Text Prompts and Auto Segmentation.
@@ 5 Conclusion
In this paper, we introduced the Segment Anything for NeRF in High Quality framework. By combining the strengths of SAM for open-world object segmentation and NeRF for aggregating information from multiple viewpoints, SANeRF-HQ represents a significant advancement in high-quality 3D segmentation. Our method was quantitatively and qualitatively evaluated on various NeRF datasets, which demonstrates SANeRF-HQ’s advantages over previous methods. In supplementary material, we demonstrate the potential of extending our work to object segmentation in 4D dynamic NeRFs. SANeRF-HQ holds promise for contributing significantly to the evolving landscape of 3D computer vision and segmentation techniques.
@@ Appendix A Implementation Details
Appendix B Efficiency Evaluation
To provide a more comprehensive understanding of the two methods storing SAM features mentioned in Section 3.1, we evaluate the efficiency of the feature distillation method and the caching method. We randomly sample three scenes from the Mip-NeRF 360 dataset as reference. By default, the pre-trained NeRF renders images at 512 $\times$ 512 as input to the SAM encoder. Under a batch size of 4,096 and a maximum iteration of 5,000, it requires on average 666.0 seconds to train the feature field for a single scene, which can then render feature maps at 64 $\times$ 64 resolution from any viewpoints at 22.4 frames per second (FPS). In contrast, the caching method can encode the images to feature maps at 3.78 FPS while using extra memory to store the SAM feature maps (64 $\times$ 64, around 4.1MB each frame). Different from encoding, the decoding process is much faster, at 168.9 FPS with the pre-computed feature maps.
@@ Appendix C Comparison with Instance Segmentation Methods
Table C.2: Comparison with Panoptic Lifting and Contrastive Lift. The results are on the data mentioned in their papers.
@@
Figure C.3: Qualitative Results of the Ray-Pair RGB Loss. The Ray-Pair RGB loss can help to recover local regions and make the results more solid.
@@ Appendix D Extending to Dynamic NeRFs
Figure D.2: Comparison with SA3D and ISRF on the Replica Room. Data is from the Others subset. SANeRF-HQ can maintain the object structure while excludes the background.
Figure D.3: Comparison with SA3D and ISRF on the Shoe Rack. Data is from the LERF subset. Our method can reproduce the segmentation details even with some occlusion.
Figure D.4: Comparison with SA3D and ISRF on Hypersim. Data is from the Others subset. ISRF contains too many false positives, while SA3D cannot cover the whole object.
Figure D.5: Comparison with SA3D and ISRF on the Espresso. Data is from the LERF subset. Our method produces the most reasonable segmentation in the distant, complex setting.
@@ Appendix E More Qualitative Results
We demonstrate the qualitative results of the Ray-Pair RGB loss in Figure C.3. The loss helps fill in the missing interior and boundaries of the masks by enforcing a local match between the similarity in labels, and the similarity in appearance.
We also provide extra qualitative comparisons between our method and other zero-shot 3D segmentation methods mentioned in the main paper. The results are given in Figures D.2D.3D.4D.5. Please watch the video for more qualitative results.
Appendix F Limitations
Though our method works well in most cases, it relies on NeRF and SAM, and its performance might be impacted by scene complexity and NeRF quality. On the other hand, the Ray-Pair RGB loss may not handle all circumstances especially given neighboring objects with identical colors and shading. Nevertheless, we present some results of our method on relatively challenging scenes to show that it may still robustly handle some of these cases, where the target objects are relatively small, in the background, partially occluded, or adjacent to other objects with similar appearance. The results are in Figure F.1 and F.2. We leave relevant potential improvements as future work.
Figure F.1: Examples of More Complex Scenarios. SANeRF-HQ can effectively segment target objects that are in the background, relatively small, and partially occluded.
Figure F.2: Examples of Objects with Similar Color. Our method can still distinguish these objects and produce reasonable results in the presence of neighbouring objects with similar appearance, where the Ray-Pair RGB loss is less helpful but remains robust.